Financial information related to In Good Taste Company for the month ended E1-13 Retained earmings statement April 30, 2010, is as follows: Obj 4 PREADSHEET Net income for April Dividends during April Retained earnings, April 1, 2010 s 125,000 18,000 395,000 Retained earnings, April 30, 2010: $502,000 Prepare a retained earnings statement for the month ended April 30, 2010.
The Role of Accounting in Business 31 E1-3 Accounting equation Ob 4 The total assets and total liabilities of Best Buy and Hewlett Packard are showi. here. Best Buy,$4,484 Best Buy (in millions) $12,758 Hewlett-Packard (in millions) $113,331 Assets Liabilities 8,274 74,389 Determine the stockholders equity of each company. E1-4 Accounting equation Obj 4 The total assets and total liabilities of Marathon Oil and Dell are shown here. Marathon Oil (in millions) $42,686 21,277 Dell (in millions) $26,500 22,229 Dell, $4,271 Assets Liabilities Determine the stockholders equity of each company. Determine the missing amount for each of the following: E1-5 Accounting equation Obi 4 Assets-Liabilities+ Stockholders Equity a. X-$35,000 + b. $80,000 c $675,000- $227,000 $60,000 $ 35,000 a.$95,000 Determine the missing amounts (in millions) for the condensed shown below. balance sheets E1-6 ccounting equation Wal-Mart Target $44,106 (b) Costco s20,682 11,490 .$30,94 Assets Liabilities Stockholders equity 98,906 64,608 13,712
Answer #1

1

Part 1

wA2hUk4hc2CTwAAAABJRU5ErkJggg==

Part 2

xxmWCepbnqwAAAAASUVORK5CYII=

Part 3

Assets =Liability + Shareholder equity

a Assets=35000+60000=$950000

b Liability=Assets- Shareholder equity=$80000-$35000=$45000

c Shareholder equity=Assets-Liability=$675000-$227000=$448000

Part 4

VoDGl8N6AAAAAElFTkSuQmCC

Assets =Liability + Shareholder equity

Target Liabilities =Assets- Shareholder equity=44106-1371=$$30394

Walmart Assets=Liability + Shareholder equity=98906+64608=$163514

Costco Shareholder equity=Assets-Liability=20682-11490=$9192